OpenStudy (anonymous):

verify green's theorem in the plane for integral (3x^2-8y^2)dx + (4y-6xy)dy, where C is the boundry of the region defined by x=0, y=0 and x + y=1

OpenStudy (turingtest):

\[\int\limits_cPdx+Qdy=\int\int\limits_D\left(\frac{\partial Q}{dx}-\frac{\partial P}{dy}\right)dA\]so first we might do well to draw D

OpenStudy (turingtest):

so it looks like our limits are pretty easy to find. Think you can tell me the bounds of our double integral?

OpenStudy (turingtest):

@kay-jay please try to stick around for a moment after posting so we can interact on this question

OpenStudy (anonymous):

am here

OpenStudy (turingtest):

cool, so do you think you can find the bounds of our double integral now that I have drawn out the region D ?

OpenStudy (anonymous):

yes but not so sure about them

OpenStudy (turingtest):

well what does x vary from?

OpenStudy (turingtest):

from what to what...?

OpenStudy (anonymous):

0 to 1

OpenStudy (turingtest):

and how about y, what is it bound by?

OpenStudy (anonymous):

y=x

OpenStudy (turingtest):

not quite

OpenStudy (turingtest):

x+y=1 so solving for y gives...?

OpenStudy (anonymous):

the question is 15 marks how about we start with line ine integration

OpenStudy (anonymous):

y=1-x

OpenStudy (turingtest):

well Green's theorem is the easier part I think doing this without greens theorem will require me to look at my notes...

OpenStudy (turingtest):

yes, so what are the bounds on y ?

OpenStudy (anonymous):

y= -1 m not so sure

OpenStudy (anonymous):

calculating partal integrals would be easy...but coming to the intervals x interval would be from 0 to 1,and y integral would be from 0 to 1-x ...(as x+y=1 is given)

OpenStudy (turingtest):

exactly^\[0\le x\le1\]\[0\le y\le1-x\]so those will be our bounds on the integral now you just need to do the\[\left(\frac{\partial Q}{dx}-\frac{\partial P}{dy}\right)\]part, which should be pretty easy by this time in the course

OpenStudy (anonymous):

then what to do now

OpenStudy (anonymous):

please help me... please

OpenStudy (turingtest):

what did you get for the expression above\[\left(\frac{\partial Q}{\partial x}-\frac{\partial P}{\partial y}\right)\]???

OpenStudy (anonymous):

i just don't get it

OpenStudy (anonymous):

oh yah let me work it out

OpenStudy (anonymous):

exectly what is the answer suppose to be

OpenStudy (turingtest):

I don't know for sure because I am having difficulty with the line integral method, but by Green's Theorem I got \(\large -\frac72\)

OpenStudy (anonymous):

ohky please do it for me and i will compare it with my solution

OpenStudy (turingtest):

I'm not going to type out the whole thing, but I will walk you through it. just do the partial derivative part\[\left(\frac{\partial Q}{\partial x}-\frac{\partial P}{\partial y}\right)\]and tell me what you get \that is a good start

OpenStudy (anonymous):

i got - \[\left(\begin{matrix}11 \\ 3\end{matrix}\right)\]

OpenStudy (anonymous):

derivative of N=-6y and the derivative of M=-16y

OpenStudy (turingtest):

ok, so one minus the other is....?

OpenStudy (anonymous):

ohk let me try it again

OpenStudy (turingtest):

\[N_x-M_y=?\]

OpenStudy (turingtest):

your partials are right, but you are just not doing the algebra correctly :P

OpenStudy (anonymous):

eix... practise makes perfect i wl kep trying

OpenStudy (anonymous):

N-M=10y

OpenStudy (turingtest):

there we go :)

OpenStudy (turingtest):

I had the wrong answer though too, let me redo it

OpenStudy (turingtest):

now I get -5/3

OpenStudy (anonymous):

i got \[\left(\begin{matrix}5 \\ 3\end{matrix}\right)\]

OpenStudy (turingtest):

oh yeah, dang negative sign :P you're right

OpenStudy (turingtest):

by the way, to write fractions write \frac53

OpenStudy (anonymous):

oh... thak you but it is 15 marks

OpenStudy (turingtest):

and put it into the equation editor, or enclose it in brackets if you know latex

OpenStudy (anonymous):

ohk thanks

OpenStudy (turingtest):

I'm fairly confident we have done the green's theorem part right, and that 5/3 is the right answer, but the line integral is giving me a headache

OpenStudy (anonymous):

and thats the part i must gain marks in cause green's theorem is only 5 marks

OpenStudy (turingtest):

I'm working on it, but I am unsure how to paramaterize t I am thinking everything from 0 to 1 let me play with it for a bit...

OpenStudy (anonymous):

please do

OpenStudy (turingtest):

so I'm thinking to parameterize the curves we do\[C_1:x=t,y=0~~~~~~~~~~~0\le t\le1\]\[C_2:x=t,y=1-t~~~~~0\le t\le1\]\[C_3:x=0,y=t~~~~~~~~~~~~0\le t\le1\]